Ida Sidha Karya Company is a family-owned company located in the village of Gianyar on the island of Bali in Indonesia. The company produces a handcrafted Balinese musical instrument called a gamelan that is similar to a xylophone. The gamelans are sold for $956. Selected data for the company
Answer #1

Hi,

Please find the detailed answer as follows:

Part A:

Unit Product Cost (Absorbtion Costing) = Direct Material + Direct Labor + Variable Overhead + Fixed Manufacturing Overhead Per Unit = 290 + 450 + 58 + 730000/19000 = $836.42

-------

Part B:

Unit Product Cost (Variable Costing) = Direct Material + Direct Labor + Variable Overhead = 290 + 450 + 58 = $798

Thanks.